OpenStudy (anonymous):

What is the slope intercept form of the equation 9x+3y=6

jimthompson5910 (jim_thompson5910):

9x+3y=6 3y=6-9x 3y=-9x+6 y=-9x/3+6/3 y=-3x+2

OpenStudy (anonymous):

y=-3x+2

OpenStudy (anonymous):

y=mx+c is slope intewrcept form...where you have m as slope and c as intercept on y axis.. y=-3x+2 slope -3 and intercept +2 ...